Two 5m x 5m thin-walled parallel plates are separated by a distance of 1 mm. Air (properties shown below at the external fluid) flows over the TOP AND BOTTOM of the parallel plate stack. You may assume fully developed hydrodynamic and thermal boundary layers between the parallel plates. Using the drawing and information below, what is the outlet mean temperature of the internal fluid?


1
Expert's answer
2021-05-31T05:51:02-0400

"\perp" Plate "\dot{Q}_{in}=m_{air}c_{pair}T_{ait}=(R*Vol)_{air}*1000*300"

"=(1.2* Area *Velocity)_{air}*3*10^{5}=1.2*25*10*3*10^{5}=9*10^7 W"

Total heat transfer to both plates "=2*\dot{Q}=198*10^7 W"

"18*10^7=\dot{m}_{H_2O}*Cp_w*(T_{out}-275)"

"18*10^7=R_{H_2O}*Vol_{H_2O}*Cp_w*(T_{out}-275)"

"18*10^7=1000*50*0.2**4118(T_{out}-275)"

"T_{out}=283.7421 K"